OpenStudy (anonymous):

A US flag measures 5ft by 7 ft. The blue portion that contains the stars measures 3 ft across the top edge and 2 ft 4 inches down the side. Compute the area of the portion of the flag containg the stripes. Help please :)

OpenStudy (gw2011):

A(US Flag)=(5)(7)=35 sq ft A(Blue Portion)=(3)(2 1/3)=7 sq ft The 2 ft 4 inches=2 4/12=2 1/3 Subtract the Blue Portion area from the Us Flag area: 35 - 7 = 28 sq ft